# DynOPETs: A Versatile Benchmark for Dynamic Object Pose Estimation and Tracking in Moving Camera Scenarios

### v1

@@ I INTRODUCTION
Object pose estimation, which involves determining an object’s 6-DoF pose (position and orientation) in 3D space, is a fundamental research challenge in computer vision and robotics. The task is crucial for various applications in embodied intelligence, augmented reality (AR), and mixed reality (MR) systems [1], such as robotic manipulation [2], dexterous hand control, and human-object interaction [3]. Despite significant advancements in recent years driven by the emergence of benchmark datasets, existing real-world datasets exhibit notable limitations. Most of the existing datasets predominantly assume static conditions - either fixed camera positions or stationary objects. This assumption contrasts sharply with real-world requirements where simultaneous camera ego-motion and object dynamics must be addressed.
The limited availability of dynamic pose estimation datasets stems primarily from annotation challenges inherent to moving-camera-dynamic-object configurations. While static-object scenarios allow for indirect pose annotation via the state-of-the-art camera pose estimation technique, such as SLAM algorithm [4] or fiducial marker-based localization method [5], dynamic settings require direct object motion tracking. In such dynamic settings, existing solutions for pose annotation rely on either labor-intensive manual annotation [6] or expensive motion capture (Mocap) systems in controlled environments [7].
@@
Another approach involves using state-of-the-art object pose estimation methods to generate high-quality pseudo-labels for annotation. However, many common objects in daily life, such as those with symmetric geometries, transparent or reflective materials, or sparse/repetitive patterns, pose significant challenges for those algorithms. As a result, a significant fraction of the object pose annotations still require manual refinement due to the low quality of the pseudo-labels.
@@
  • We conduct a systematic benchmark of various taxonomy-specific pose estimation and object pose tracking methods tailored to the unique characteristics of our dataset.
@@
CAMERA25 [8] 6 1085 - 300K 4M
Omni6D [9] 166 4688 - 0.8M 5.6M
SOPE [10] 149 4162 - 475K 5M
YCB-Video [11] - 21 92 133K 0.6M
T-LESS [12] - 30 20 48K 48K
Linemod [13] - 15 - 18K 15k
StereoOBJ-1M [14] - 18 182 393K 1.5M
YCBInEOAT [6] - 5 9 7449 3.7K
HOT3D [7] - 33 425 3.7M 6M
@@
REAL275 [8] 6 42 18 8K 35K
Wild6D [15] 5 162 486 10K 10K
PhoCaL [16] 8 60 24 3.9K 91K
HANDAL [17] 17 212 2.2K 308K 1.2M
HouseCat6D [18] 10 194 41 23.5K 160K
ROPE [10] 149 581 363 332K 1.5M
@@ II RELATED WORK
According to  [19], We categorize object pose estimation methods into 4 classes: instance-level object pose estimation (IOPE), unseen object pose estimation (UOPE), category-level object pose estimation (COPE), and object pose tracking (OPT). This section discusses representative and state-of-the-art methods within each class, followed by an overview of recently published datasets relevant to these areas.
@@ II-A1 Instance-level Object Pose Estimation
These methods predict the pose of only those objects present in the training set, often relying on priors such as CAD models or posed reference views during inference [20]. Object pose estimation can be achieved through template matching [21] or feature correspondence matching [22] given these priors. Some approaches directly predict object pose using neural networks [20], while others employ voting schemes [23]. Instance-level methods are the most mature but also the most restrictive due to their inability to generalize to novel objects. Consequently, research has shifted towards the more generalizable UOPE and COPE methods, which are the primary focus of this paper and our proposed dataset.
@@ II-A2 Unseen Object Pose Estimation
Similar to IOPE methods, UOPE methods rely on CAD models or posed reference views as priors. However, UOPE methods can generalize to novel objects, enabling pose prediction for unseen objects during inference. Recent works in this area are broadly categorized into two main strategies: feature matching and template matching-based methods. Feature matching methods [24, 25, 26, 27] extract robust image features and establish correspondences with generic, object-agnostic CAD model features. These approaches enable zero-shot estimation but may rely heavily on feature discriminability. In contrast, template matching methods [28, 29, 30, 1] leverage extensive sets of object templates to achieve strong generalization across novel objects.
@@ II-A3 Category-level Object Pose Estimation
COPE methods typically handle novel objects within the same category by learning category-specific shape representations, eliminating the reliance on CAD models and reference views. Wang et al. [8] pioneered this approach by introducing the normalized object coordinate space (NOCS), a canonical representation where all objects within the same category are consistently aligned in terms of size and orientation. Aligning predicted canonical representations with depth points allows for the estimation of object pose.
Recent advancements have explored different network designs, such as using vision transformers [31] or diffusion models [32] to predict better canonical representations. Several works [33, 34, 35, 32, 36] employ deformed categorical shape templates to improve canonical representations, while others [37, 10] employ diffusion models to formulate COPE as a conditional generative task. Furthermore, some methods directly predict object pose, employing either enhanced geometric feature extractors [38] or a combination of geometric and semantic priors [39, 40, 41].
@@ II-A4 Object Pose Tracking
Unlike single-frame object pose estimation methods, object pose tracking leverage temporal information across video sequences to estimate object poses [19]. Typically, tracking methods also require posed reference views or CAD models. While classical methods [42] often track object poses using geometric or feature-based approaches, recent methods [43] jointly track and refine object shape through neural implicit representations, providing robust performance. Moreover, Wen et al. [1] leverage large-scale pre-trained models and efficiently track unseen objects using large-scale synthetic training data for enhanced generalization.
@@ II-B Object Pose Estimation Datasets
Linemod [13] introduced a standard instance-level RGB-D benchmark featuring texture-less objects with limited occlusions in relatively simple scenarios. YCB-Video [11] subsequently expanded realism with household objects. T-Less [12] introduced texture-less but symmetric objects in challenging industrial scenarios, highlighting pose ambiguity issues. REAL275 [8] established the first real-world benchmark for category-level pose estimation by introducing canonical object representations across 6 categories. StereOBJ-1M [14] further tackled challenging materials with a much larger scale, such as transparent and reflective objects. To further diversify modalities, PhoCaL [16] provided multimodal images focusing on challenging transparent and reflective household items with 8 categories. Wild6D [15] advanced realism further by capturing large-scale videos of objects in the wild. More recently, HANDAL [17] focused on graspable objects, collecting 17 categories of manipulable items in cluttered scenes via multi-view imagery and annotating poses with [43]. HOT3D [3] was the first to address egocentric scenarios featuring dense hand–object interactions. However, it incorporates a limited variety of objects and constrained scene settings. HouseCat6D [18] delivered highly accurate annotations for 10 categories of household objects using polarization imagery to enhance robustness to reflections. OmniNOCS [31] was a unified NOCS dataset that integrated data from multiple domains and featured over 90 object classes, making it the largest NOCS dataset to date. Omni6D [9] and ROPE [10] also provided a large-scale dataset for object pose estimation.
Figure 2: Overview. The left block is CAD model scanning and preprocessing, as well as hand-eye calibration. The right block is the data acquisition and efficient object pose annotation pipeline. The dynamic object is captured by a moving RGB-D camera, whose ground truth pose annotation is calculated by fusing refined absolute object pose estimates and relative object pose estimates. The camera pose is provided by a Mocap.
@@ III-B Hardware Setup and Objects
RGB-D data was captured using a consumer-grade iPad Pro with the hardware-synchronized Structure Sensor Pro [44] depth camera. CAD models for all 175 objects were scanned using the sensors’ integrated RGB-D 3D modelling software. Following mesh reconstruction, objects were manually aligned so that their centers coincided with the origin of the 3D coordinate system, and their front, top, and side views aligned with the x, y, and z axes, respectively. Consistent with the definition in Wang et al. [8], the coordinate system of an object’s CAD model is defined as Object Local Coordinates (OLC) system, also known as the canonical space in the COPE domain, as illustrated in Figure 3. Additionally, we employed SAM2 [45] to automatically segment objects in these frames.
@@ III-C Sensor Calibration and Synchronization
Intrinsic and extrinsic camera parameters were calibrated using the depth camera SDK, which provides simultaneous calibration of both RGB and depth cameras. We developed a custom camera driver based on the SDK that warps the depth map to the RGB camera’s coordinate frame, resulting in co-registered RGB-D video sequences. A Mocap system provided camera poses. The relative transformation between the Mocap marker frame and the camera’s optical center was calibrated using a hand-eye calibration procedure, yielding camera poses within the Mocap coordinate system. A temporal offset existed between the Mocap system’s timestamps and the RGB-D frame timestamps. To resolve this, we used visual odometry (VO) [4] to estimate the camera pose for each frame and registered these VO-derived poses to the Mocap poses. This registration process established a mapping between each camera frame and the corresponding Mocap timestamp.
@@ III-D Efficient Object Pose Annotation
To provide context for our approach, we begin with key definitions. Firstly, the absolute object pose, which we aim to annotate, is defined as the Sim(3) transformation (scale, rotation, and translation) required to align an object from its OLC system to the camera frame. Since our CAD models and corresponding point clouds are scaled to real-world dimensions, this Sim(3) simplifies to an SE(3) transformation (rotation and translation). Secondly, the relative object pose describes the SE(3) transformation between the object’s poses at two different time steps in the camera coordinate system. These two definitions are directly relevant to the following two approaches for obtaining the object pose annotations.
@@ III-D1 Absolute Object Pose Estimation and Refinement
The first approach involves directly generating high-quality pseudo pose annotations using existing state-of-the-art algorithms. Model-based method FoundationPose [1] stands out in this category, leveraging CAD models to output object motion. Although generally accurate, FoundationPose [1] can occasionally generate discontinuous pose estimates. To address this issue, we adopt a global EKF [46] with Rauch-Tung-Striebel (RTS) smoother [47] to refine the object pose predictions, mitigating pose jitters. The RTS smoother employs a constant velocity model for both translation $p\in\mathbf{R}^{3}$ and rotational quaternion $q\in\mathbb{S}^{3}$, where the state variables are defined as $\mathbf{x}=[p,q,\upsilon,\omega]$, where $\upsilon\in\mathbb{R}^{3},\omega\in\mathbb{R}^{3}$ refers the velocity and angular velocity. We first perform a forward pass eq. 2 using the standard EKF prediction and update equations to obtain filtered state estimates $\mathbf{x}_{t}$ from step $t-1$ to step $t$, then, a backward pass eq. 3 using the RTS smoother refines these estimates from step $t+1$ to $t$:
@@
where $\mathbf{z}$ refers to the observed pose, and $\mathbf{F}$ and $\mathbf{H}$ represent the state transition and observation matrices, respectively. $K$ is the extended Kalman gain used in the update step [46], $\mathbf{P}$ denotes the covariance matrix, and $\mathbf{G}$ represents the smoothing gain. After our refinement, the pose annotations of objects with rich textures and distinctive shapes are now accurate and smooth, while for those objects exhibiting significant symmetry or lacking distinctive texture, such as objects within the bowl category, the pose annotation is still not high-quality enough.
@@ III-D2 Relative Object Pose Estimation via Point Tracking
BundleTrack [42] serves as a representative example of this approach, employing keypoint detection and matching to establish 2D correspondences. These 2D correspondences are then back-projected into 3D space using sensor depths, forming 3D-3D correspondences that enable the computation of relative object poses via point registration techniques. Our method adopts a similar strategy but replaces the keypoint detection and matching component with a state-of-the-art Track Any Point (TAP) model [48]. TAP models accept a video sequence as input and output 2D correspondence points for a group of specified query points in the reference frame across all frames. Compared to traditional feature extraction and matching methods, TAP model [48] is better equipped to leverage the temporal information inherent in video sequences, enabling them to effectively handle objects with limited texture. While this method produces smoother object pose estimates and is more robust to symmetry ambiguities, it is susceptible to cumulative errors, leading to drift in the estimated trajectory over time.
@@ III-D3 Object Pose Graph Optimization
NOCS [8] 73.50 48.75 20.47 11.57 16.71 22.43 32.98
NOCS (MV-ROPE)* [49] 88.74 71.67 38.19 28.84 31.94 48.50 55.53
DiffusionNOCS [32] 83.00 38.89 4.57 39.87 43.83 55.82 64.80
VI-Net [39] 95.66 87.34 52.35 55.09 59.98 72.35 79.70
IST-Net [35] 97.52 90.68 58.99 51.57 52.85 70.90 73.41
GenPose [37] - - - 47.48 54.30 60.96 71.84
SecondPose [40] 98.55 90.90 53.83 58.57 64.00 75.13 83.29
AGPose [41] 98.47 92.94 66.34 61.61 63.24 80.02 82.06
@@
SecondPose [40] 98.53 92.24 57.42 61.88 67.05 78.24 84.08
AGPose [41] 98.52 92.10 68.53 67.11 67.50 81.23 81.92
SecondPose* [40] 99.38(+0.85) 96.23(+3.99) 59.80(+2.38) 76.33(+14.45) 82.91(+15.86) 89.43(+11.19) 97.19(+13.11)
AGPose* [41] 99.91(+1.39) 99.76(+7.66) 84.42(+15.89) 75.10(+8.99) 75.74(+8.24) 94.35(+13.12) 96.46(+14.54)

### v2

# IEEEexample.bib

@@ I INTRODUCTION
Object pose estimation, which involves determining an object’s 6-DoF pose (position and orientation) in 3D space, is a fundamental research challenge in computer vision and robotics. The task is crucial for various applications in embodied intelligence, augmented reality (AR), and mixed reality (MR) systems [wen2024foundationpose], such as robotic manipulation [ubellacker2024high], dexterous hand control, and human-object interaction [banerjee2025hot3d]. Despite significant advancements in recent years driven by the emergence of benchmark datasets, existing real-world datasets exhibit notable limitations. Most of the existing datasets predominantly assume static conditions - either fixed camera positions or stationary objects. This assumption contrasts sharply with real-world requirements where simultaneous camera ego-motion and object dynamics must be addressed.
The limited availability of dynamic pose estimation datasets stems primarily from annotation challenges inherent to moving-camera-dynamic-object configurations. While static-object scenarios allow for indirect pose annotation via the state-of-the-art camera pose estimation technique, such as SLAM algorithm [teed2021droid] or fiducial marker-based localization method [olson2011apriltag], dynamic settings require direct object motion tracking. In such dynamic settings, existing solutions for pose annotation rely on either labor-intensive manual annotation [wen2020se] or expensive motion capture (Mocap) systems in controlled environments [drost2017introducing].
@@
Another approach [guo2023handal] involves using state-of-the-art object pose estimation methods [wen2024foundationpose] to generate high-quality pseudo-labels for annotation. However, many common objects in daily life, such as those with symmetric geometries, transparent or reflective materials, or sparse/repetitive patterns, pose significant challenges for those algorithms. As a result, a significant fraction of the object pose annotations still require manual refinement due to the low quality of the pseudo-labels.
@@
  • We conduct a systematic benchmark of various taxonomy-specific pose estimation and object pose tracking methods on our dataset.
@@
CAMERA25 [wang2019normalized] 6 1085 - 300K 4M
Omni6D [zhang2024omni6d] 166 4688 - 0.8M 5.6M
SOPE [zhang2024omni6dpose] 149 4162 - 475K 5M
YCB-Video [xiang2018posecnn] - 21 92 133K 0.6M
T-LESS [hodan2017t] - 30 20 48K 48K
Linemod [hinterstoisser2013model] - 15 - 18K 15k
StereoOBJ-1M [liu2021stereobj] - 18 182 393K 1.5M
YCBInEOAT [wen2020se] - 5 9 7449 3.7K
HOT3D [banerjee2025hot3d] - 33 425 3.7M 6M
@@
REAL275 [wang2019normalized] 6 42 18 8K 35K
Wild6D [fu2022category] 5 162 486 10K 10K
PhoCaL [wang2022phocal] 8 60 24 3.9K 91K
HANDAL [guo2023handal] 17 212 2.2K 308K 1.2M
HouseCat6D [jung2024housecat6d] 10 194 41 23.5K 160K
ROPE [zhang2024omni6dpose] 149 581 363 332K 1.5M
@@ II RELATED WORK
According to  [liu2024deep], We categorize object pose estimation methods into 4 classes: instance-level object pose estimation (IOPE), unseen object pose estimation (UOPE), category-level object pose estimation (COPE), and object pose tracking (OPT). This section discusses representative and state-of-the-art methods within each class, followed by an overview of recently published datasets relevant to these areas.
@@ II-A1 Instance-level Object Pose Estimation
These methods predict the pose of only those objects present in the training set, often relying on priors such as CAD models or posed reference views during inference [wang2019densefusion]. Object pose estimation can be achieved through template matching [dang2022learning] or feature correspondence matching [xu2022rnnpose] given these priors. Some approaches directly predict object pose using neural networks [wang2019densefusion], while others employ voting schemes [zhou2023deep]. Instance-level methods are the most mature but also the most restrictive due to their inability to generalize to novel objects. Consequently, research has shifted towards the more generalizable UOPE and COPE methods, which are the primary focus of this paper and our proposed dataset.
@@ II-A2 Unseen Object Pose Estimation
Similar to IOPE methods, UOPE methods rely on CAD models or posed reference views as priors. However, UOPE methods can generalize to novel objects, enabling pose prediction for unseen objects during inference. Recent works in this area are broadly categorized into two main strategies: feature matching and template matching-based methods. Feature matching methods [chen2023zeropose, lin2024sam, caraffa2024freeze, huang2024matchu] extract robust image features and establish correspondences with generic, object-agnostic CAD model features. These approaches enable zero-shot estimation but may rely heavily on feature discriminability. In contrast, template matching methods [labbe2022megapose, nguyen2024gigapose, ornek2024foundpose, wen2024foundationpose] leverage extensive sets of object templates to achieve strong generalization across novel objects.
@@ II-A3 Category-level Object Pose Estimation
COPE methods typically handle novel objects within the same category by learning category-specific shape representations, eliminating the reliance on CAD models and reference views. Wang et al. [wang2019normalized] pioneered this approach by introducing the normalized object coordinate space (NOCS), a canonical representation where all objects within the same category are consistently aligned in terms of size and orientation. Aligning predicted canonical representations with depth points allows for the estimation of object pose.
Recent advancements have explored different network designs, such as using vision transformers [krishnan2024omninocs] or diffusion models [ikeda2024diffusionnocs] to predict better canonical representations. Several works [wang2023query6dof, meng2023kgnet, liu2023net, ikeda2024diffusionnocs, wang2024gs] employ deformed categorical shape templates to improve canonical representations, while others [li2023generative, zhang2024omni6dpose] employ diffusion models to formulate COPE as a conditional generative task. Furthermore, some methods directly predict object pose, employing either enhanced geometric feature extractors [li2025gce] or a combination of geometric and semantic priors [lin2023vi, chen2024secondpose, lin2024instance].
@@ II-A4 Object Pose Tracking
Unlike single-frame object pose estimation methods, object pose tracking leverage temporal information across video sequences to estimate object poses [liu2024deep]. Typically, tracking methods also require posed reference views or CAD models. While classical methods [wen2021bundletrack] often track object poses using geometric or feature-based approaches, recent methods [wen2023bundlesdf] jointly track and refine object shape through neural implicit representations, providing robust performance. Moreover, Wen et al. [wen2024foundationpose] leverage large-scale pre-trained models and efficiently track unseen objects using large-scale synthetic training data for enhanced generalization.
@@ II-B Object Pose Estimation Datasets
Linemod [hinterstoisser2013model] introduced a standard instance-level RGB-D benchmark featuring texture-less objects with limited occlusions in relatively simple scenarios. YCB-Video [xiang2018posecnn] subsequently expanded realism with household objects. T-Less [hodan2017t] introduced texture-less but symmetric objects in challenging industrial scenarios, highlighting pose ambiguity issues. REAL275 [wang2019normalized] established the first real-world benchmark for category-level pose estimation by introducing canonical object representations across 6 categories. StereOBJ-1M [liu2021stereobj] further tackled challenging materials with a much larger scale, such as transparent and reflective objects. To further diversify modalities, PhoCaL [wang2022phocal] provided multimodal images focusing on challenging transparent and reflective household items with 8 categories. Wild6D [fu2022category] advanced realism further by capturing large-scale videos of objects in the wild. More recently, HANDAL [guo2023handal] focused on graspable objects, collecting 17 categories of manipulable items in cluttered scenes via multi-view imagery and annotating poses with [wen2023bundlesdf]. HOT3D [banerjee2025hot3d] was the first to address egocentric scenarios featuring dense hand–object interactions. However, it incorporates a limited variety of objects and constrained scene settings. HouseCat6D [jung2024housecat6d] delivered highly accurate annotations for 10 categories of household objects using polarization imagery to enhance robustness to reflections. OmniNOCS [krishnan2024omninocs] is a unified NOCS dataset that integrated data from multiple domains and featured over 90 object classes, making it the largest NOCS dataset to date. Omni6D [zhang2024omni6d] and ROPE [zhang2024omni6dpose] also provided a large-scale dataset for object pose estimation.
Figure 2: Overview. The left block is CAD model scanning and preprocessing, as well as hand-eye calibration. The right block is the data acquisition and efficient object pose annotation pipeline. The dynamic object is captured by a moving RGB-D camera, and the ground truth object pose annotation is calculated by fusing refined absolute object pose estimates and relative object pose estimates. The camera pose is provided by a Mocap.
@@ III-B Hardware Setup and Objects
RGB-D data was captured using a consumer-grade iPad Pro with the hardware-synchronized Structure Sensor Pro [structuresensor] depth camera. CAD models for all 175 objects were scanned using the sensors’ integrated RGB-D 3D modelling software. Following mesh reconstruction, objects were manually aligned so that their centers coincided with the origin of the 3D coordinate system, and their front, top, and side views aligned with the x, y, and z axes, respectively. Consistent with the definition in Wang et al. [wang2019normalized], the coordinate system of an object’s CAD model is defined as Object Local Coordinates (OLC) system, also known as the canonical space in the COPE domain, as illustrated in Figure 3. Additionally, we employed SAM2 [ravi2024sam] to automatically segment objects in these frames.
@@ III-C Sensor Calibration and Synchronization
Intrinsic and extrinsic camera parameters were calibrated using the depth camera SDK, which provides simultaneous calibration of both RGB and depth cameras. We developed a custom camera driver based on the SDK that warps the depth map to the RGB camera’s coordinate frame, resulting in co-registered RGB-D video sequences. A Mocap system provided camera poses. The relative transformation between the Mocap marker frame and the camera’s optical center was calibrated using a hand-eye calibration procedure, yielding camera poses within the Mocap coordinate system. A temporal offset existed between the Mocap system’s timestamps and the RGB-D frame timestamps. To resolve this, we used visual odometry (VO) [teed2021droid] to estimate the camera pose for each frame and registered these VO-derived poses to the Mocap poses. This registration process established a mapping between each camera frame and the corresponding Mocap timestamp.
@@ III-D Efficient Object Pose Annotation
Illustrated in fig. 2, To provide context for our approach, we begin with key definitions. Firstly, the absolute object pose, which we aim to annotate, is defined as the Sim(3) transformation (scale, rotation, and translation) required to align an object from its OLC system to the camera frame. Since our CAD models and corresponding point clouds are scaled to real-world dimensions, this Sim(3) simplifies to an SE(3) transformation (rotation and translation). Secondly, the relative object pose describes the SE(3) transformation between the object’s poses at two different time steps in the camera coordinate system. These two definitions are directly relevant to the following two approaches for obtaining the object pose annotations.
@@ III-D1 Absolute Object Pose Estimation and Refinement
The first approach involves directly generating high-quality pseudo pose annotations using existing state-of-the-art algorithms. Model-based method FoundationPose [wen2024foundationpose] stands out in this category, leveraging CAD models to output object motion. Although generally accurate, FoundationPose [wen2024foundationpose] can occasionally generate discontinuous pose estimates. To address this issue, we adopt a global EKF [bishop2001introduction] with Rauch-Tung-Striebel (RTS) smoother [rauch1965maximum] to refine the object pose predictions, mitigating pose jitters. The RTS smoother employs a constant velocity model for both translation $p\in\mathbf{R}^{3}$ and rotational quaternion $q\in\mathbb{S}^{3}$, where the state variables are defined as $\mathbf{x}=[p,q,\upsilon,\omega]$, where $\upsilon\in\mathbb{R}^{3},\omega\in\mathbb{R}^{3}$ refers the velocity and angular velocity. We first perform a forward pass eq. 2 using the standard EKF prediction and update equations to obtain filtered state estimates $\mathbf{x}_{t}$ from step $t-1$ to step $t$, then, a backward pass eq. 3 using the RTS smoother refines these estimates from step $t+1$ to $t$:
@@
where $\mathbf{z}$ refers to the observed pose, and $\mathbf{F}$ and $\mathbf{H}$ represent the state transition and observation matrices, respectively. $K$ is the extended Kalman gain used in the update step [bishop2001introduction], $\mathbf{P}$ denotes the covariance matrix, and $\mathbf{G}$ represents the smoothing gain. After our refinement, the pose annotations of objects with rich textures and distinctive shapes are now accurate and smooth, while for those objects exhibiting significant symmetry or lacking distinctive texture, such as objects within the bowl category, the pose annotation is still not high-quality enough.
@@ III-D2 Relative Object Pose Estimation via Point Tracking
BundleTrack [wen2021bundletrack] serves as a representative example of this approach, employing keypoint detection and matching to establish 2D correspondences. These 2D correspondences are then back-projected into 3D space using sensor depths, forming 3D-3D correspondences that enable the computation of relative object poses via point registration techniques. Our method adopts a similar strategy but replaces the keypoint detection and matching component with a state-of-the-art Track Any Point (TAP) model [karaev2024cotracker3]. TAP models accept a video sequence as input and output 2D correspondence points for a group of specified query points in the reference frame across all frames. Compared to traditional feature extraction and matching methods, TAP model [karaev2024cotracker3] is better equipped to leverage the temporal information inherent in video sequences, enabling them to effectively handle objects with limited texture. While this method produces smoother object pose estimates and is more robust to symmetry ambiguities, it is susceptible to cumulative errors, leading to drift in the estimated trajectory over time.
@@ III-D3 Object Pose Graph Optimization
NOCS [wang2019normalized] 73.50 48.75 20.47 11.57 16.71 22.43 32.98
NOCS (MV-ROPE)* [yang2024mv] 88.74 71.67 38.19 28.84 31.94 48.50 55.53
DiffusionNOCS [ikeda2024diffusionnocs] 83.00 38.89 4.57 39.87 43.83 55.82 64.80
GCASP [li2023generative] 89.13 52.84 14.79 50.22 52.74 65.89 70.97
VI-Net [lin2023vi] 95.66 87.34 52.35 55.09 59.98 72.35 79.70
IST-Net [liu2023net] 97.52 90.68 58.99 51.57 52.85 70.90 73.41
GenPose [zhang2023generative] - - - 47.48 54.30 60.96 71.84
SecondPose [chen2024secondpose] 98.55 90.90 53.83 58.57 64.00 75.13 83.29
AGPose [lin2024instance] 98.47 92.94 66.34 61.61 63.24 80.02 82.06
@@
SecondPose [chen2024secondpose] 98.53 92.24 57.42 61.88 67.05 78.24 84.08
AGPose [lin2024instance] 98.52 92.10 68.53 67.11 67.50 81.23 81.92
SecondPose* [chen2024secondpose] 99.38(+0.85) 96.23(+3.99) 59.80(+2.38) 76.33(+14.45) 82.91(+15.86) 89.43(+11.19) 97.19(+13.11)
AGPose* [lin2024instance] 99.91(+1.39) 99.76(+7.66) 84.42(+15.89) 75.10(+8.99) 75.74(+8.24) 94.35(+13.12) 96.46(+14.54)